You buy a 20-year bond with a coupon rate of 7.6% that has a yield to maturity of 9.7%. (Assume a face value of $1,000 and semiannual coupon payments.) Six months later, the yield to maturity is 10.7%. What is your return over the 6 months? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Negative amount should be indicated by a minus sign.)

Rate of return = %
Answer #1

Solution :-

bond today Step 1+ Callulate PV ot $1000 FV int do matndy 20 years Team Time to maturity 4o Stmi annual periods Semi - annual1000 6.64811003 38 1- 6.64871002 O.0485 S64811002 6.64871003 1000 38 0.0485 6.64871007 665. 66 +1S0. 41 $816.07 Step 2 Calulapv atter Six months FV + -14 1 38 0.o535 1000 + 39 l+0.0535) 39 I+0.0535) 38 O.OS35 1000 + .O535)39 loS35)39 1000 7.63391039

Step 3 Calculate Rodte of Retuan Rate ot Return C+(Py-Po 3.8 t (748.23-816.01) 816.07 38 t(67.84) 38-67.84 816.07 -29. 84 816

Step 1 - Calculate PV of bond today & FV = $1000 Tint to matuonty (t) = 20 years or no Term = 20 years Time to maturity = 40 semi-annual A periods Semi - annual coupon = $1000 x 3,8y. (0) - $38 semi-annual Yield ly2 – 9.7% – 4.85% or 0.0485 Now, PV = { 1 - 2 423 4] + Avt - 2042 [1 (14.09esto toho could go o slo (10489)+

= yes 1-6,64811003) + ca. 249 100 38 (0.0485 x 5.64871002 6.64871002) 16.64871002 = - 665. 66 t 150.41 $816.07 PV of bond after Step 2 - Calculate six months Remaining steam = 20 years - 0.5 years I 19.5 years (t) Semi-annual periods = 19.5x2 – 39 Vield to matusity (Y) = 10.7% (Semi-annual) a 5.35% or 0.0535

PV after six months 38 0.0535 L 1) + 1000 +0.053539 0.0535 L 1000 (1.053539 35) = 2053511- 7.6339)039) + 7.63691039 -1000 7 7.63391039 = - 6. 63391039 0.0535 7. 63391039) 617,24 + 130.99 $748.23

Step 3 – Calculate Rate of Return Rate of Return - C+(PV, - PV.) - PVO - 38+ (748.23 - 816.07) 816.07 - 38+ (-67.84) 816.07 38 - 67.84 816.07 = - -29.84 816.07 -0.0366 or 3.66% Where, Ca Semi-annual Coupon PU = PV after 6 months PV = PU to day